What is the length of the transverse axis of the hyperbola defined by the equation below? (y-6)^2/6^2-(x+5)^2/12^2=1 Answer by lwsshak3(11628) 
 What is the length of the transverse axis of the hyperbola defined by the equation below
(y-6)^2/6^2-(x+5)^2/12^2=1
..
Standard form of hyperbola with horizontal transverse axis: (x-h)^2/a^2-(y-k)^2/b^2=1,
with (h,k) being the (x,y) coordinates of the center.
Standard form of hyperbola with vertical transverse axis: (y-k)^2/a^2-(x-h)^2/b^2=1, with (h,k) being the (x,y) coordinates of the center.
The difference between these two forms is that the (x-h)^2 and (y-k)^2 terms are interchanged.
..
Given equation of hyperbola: (y-6)^2/6^2-(x+5)^2/12^2=1
This is a hyperbola of the 2nd form listed. It has a vertical transverse axis with center at (-5,6).
a^2=6^2
a=6=distance from center to vertex
Length of transverse axis=distance between vertices=2a=12
See the graph below for visual evidence of the algebra above:
..
y=(36(1+(x+5)^2/144))^.5+6